OpenStudy (madgirlwithabluebox):

Will fan and medal. solve the equation. check the solution. -2/x+4=4/x+3

OpenStudy (shaik0124):

-2(x+3)=4(x+4)

OpenStudy (shaik0124):

can u solve now

OpenStudy (madgirlwithabluebox):

-2x-6=4x+16?

OpenStudy (shaik0124):

now make similar terms on one side

OpenStudy (shaik0124):

-2x-4x=16+6

OpenStudy (madgirlwithabluebox):

-16-6=4x+2x

OpenStudy (madgirlwithabluebox):

-22=6x?

OpenStudy (madgirlwithabluebox):

-22/6=6x/6

OpenStudy (madgirlwithabluebox):

-11/3?

OpenStudy (shaik0124):

yes -11/3 is the answer

OpenStudy (shaik0124):

u got it